OpenStudy (anonymous):

Which of the following points is in the solution set of y < x2 - 2x - 8? (-2, -1) (0, -2) (4, 0)

OpenStudy (anonymous):

(4,-2)

jhonyy9 (jhonyy9):

solve the quadratic equation x^2 -2x -8=0 (x-4)(x+2)=0 x-4=0 x=4 x+2=0 x=-2 - so from this result that 4 and -2 are roots of this equation - because there is x^2 -2x -8 >y so this signing ,mean that .... so - sorry - answer on this question i think that you can getting just substituting these values of x and check what values of y will satisfie this property that x^2 -2x -8 is greater than y hope that is understandably good luck bye

OpenStudy (anonymous):

thx